OpenStudy (anonymous):

For the function f(x) = –2(x + 3)2 − 1, identify the vertex, domain, and range.

OpenStudy (anonymous):

use desmos.com

OpenStudy (solomonzelman):

f(x) is a polynomial function (why? cuz it has whole number powers of x only, and no other complex terms), and a polynomial is always continuous on the interval (-∞,+∞). that is for your domain.

OpenStudy (solomonzelman):

the vertex of a parabola (your function is a parabola) in a form of: f(x)=a(x-h)²+k is the point: (h,k)

OpenStudy (solomonzelman):

can you identify the vertex here? \(f(x)=-2(x+3)^2-1\)

OpenStudy (anonymous):

Is it (3, -1)?

OpenStudy (solomonzelman):

almost.... \(f(x)=-2(x+3)^2-1\) \(f(x)=-2(x-\color{red}{-3})^2-\color{red}{1}\)

OpenStudy (solomonzelman):

it is (-3,1)

OpenStudy (solomonzelman):

and the domain is all real numbers (as I have reviously explained)

OpenStudy (solomonzelman):

Now, your leading coefficient is negative, that means your parabola opens down and goes into -∞..... BUT, the range will be limited, and it is limited by the vertex because vertex is the maximum point (in any case when parabola opens down)

OpenStudy (solomonzelman):

what is the y-value of your vertex? that is where the range ends.... and it starts from negative infinity

OpenStudy (anonymous):

Wouldn't the range be y is equal to or lesser than -3

OpenStudy (solomonzelman):

-3 is the x-coordinate of the vertex....

OpenStudy (solomonzelman):

so not to -3, but to?

OpenStudy (anonymous):

This it is -1

OpenStudy (solomonzelman):

yes, the range is y≤-3

OpenStudy (solomonzelman):

oh sorry

OpenStudy (solomonzelman):

I mean y≤1

OpenStudy (anonymous):

ok thank you

OpenStudy (solomonzelman):

yw